OpenStudy (anonymous):

Okay, Can someone please help me and show me the steps to show the prob. Thanks! Here is the problem: Tom has 117 Quarters and dimes worth $15.75. How many of each coin does he have, btw this is a linear equation....

OpenStudy (anonymous):

x(.25)+y(.1)=15.75 x+y=117 Try solving this system of equations

OpenStudy (anonymous):

umm ok.....

OpenStudy (anonymous):

let me know if you are stuck

OpenStudy (anonymous):

i will notify you if i need help

OpenStudy (anonymous):

1st step x(.25) + y(.1) = 15.75 (x + y)(-.25) = (117) (-.25) or x(.25) + y(.1) + 15.75 (x + y)(-.1) = (117) (-.1)

OpenStudy (anonymous):

that kinda makes not sense

OpenStudy (anonymous):

forget it! Imma ask my math teahcer